# Slow Transition to Low-Dimensional Chaos in Heavy-Tailed Recurrent Neural Networks

### v1

@@ Abstract
Growing evidence suggests that synaptic weights in the brain follow heavy-tailed distributions, yet most theoretical analyses of recurrent neural networks (RNNs) assume Gaussian connectivity. We systematically study the activity of RNNs with random weights drawn from biologically plausible Lévy alpha-stable distributions. While mean-field theory for the infinite system predicts that the quiescent state is always unstable—implying ubiquitous chaos—our finite-size analysis reveals a sharp transition between quiescent and chaotic dynamics. We theoretically predict the gain at which the system transitions from quiescent to chaotic dynamics, and validate it through simulations. Compared to Gaussian networks, heavy-tailed RNNs exhibit a broader parameter regime near the edge of chaos, namely a slow transition to chaos. However, this robustness comes with a tradeoff: heavier tails reduce the Lyapunov dimension of the attractor, indicating lower effective dimensionality. Our results reveal a biologically aligned tradeoff between the robustness of dynamics near the edge of chaos and the richness of high-dimensional neural activity. By analytically characterizing the transition point in finite-size networks—where mean-field theory breaks down—we provide a tractable framework for understanding dynamics in realistically sized, heavy-tailed neural circuits.
@@ 1 Introduction
Advances in connectomics yield increasingly detailed wiring diagrams of neural circuits across species and brain regions (Dorkenwald et al., 2024; The MICrONS Consortium, 2025). This progress raises fundamental questions: what structural principles govern neural circuits, and how do they support the brain’s remarkable computational power? One such prominent structural feature is the presence of heavy-tailed Foss et al. (2011) synaptic weight distributions, consistently observed across the mammalian cortex (Song et al., 2005; Lefort et al., 2009; Dorkenwald et al., 2022), mammalian hippocampus (Ikegaya et al., 2013), and even in the Drosophila central brain (Scheffer et al., 2020). Notably, this feature stands in sharp contrast to the Gaussian weight assumptions that dominate theoretical neuroscience studies and light-tailed distributions utilized in the standard initialization schemes in modern artificial neural networks LeCun et al. (2002); Glorot and Bengio (2010); He et al. (2015). One way to formally model heavy tails is with the family of Lévy $\alpha$-stable distributions (Feller, 1971; Borak et al., 2005), which emerges as a natural generalization of the familiar Gaussian distribution ($\alpha=2$) via the generalized central limit theorem. The family is parameterized by a stability index $\alpha$, where smaller values of $\alpha$ correspond to heavier tails. For $\alpha<2$, these distributions feature heavy, power-law tails. Similarly to experimentally measured synaptic weights, samples generated from such distributions consistently contain large outliers that dominate many sample statistics. As we show in this paper, this can strongly affect neural dynamics.
@@
One key phenomenon studied in theoretical neuroscience and machine learning is the transition to chaos, which has long been hypothesized to support optimal information flow and computational capacity at the so-called edge of chaos in a wide variety of randomly initialized neural networks. This encompasses RNNs (Bertschinger et al., 2004; Legenstein and Maass, 2007; Toyoizumi and Abbott, 2011; Schuecker et al., 2018) and feedforward neural networks (Schoenholz et al., 2016; Poole et al., 2016). In the context of feedforward networks, this effect can be understood in terms of the neural network Gaussian process kernel (Neal, 1996): outside of the critical point, analogous to the edge of chaos in RNNs, a trivial fixed point with kernel constant almost everywhere is approached exponentially fast, limiting the effective depth of information propagation and network trainability (Schoenholz et al., 2016; Lee et al., 2017). The transition between chaotic and non-chaotic dynamics in RNNs is often discussed in terms of eigenvalues of the weight matrix (Rajan and Abbott, 2006; Aljadeff et al., 2015). According to the circular law (Girko, 1985; Tao et al., 2010), its eigenvalues are bounded in a circle of radius proportional to the standard deviation of the distribution of weight entries. If the standard deviation is large enough, some eigenvalues fall outside of the unit circle and the quiescent state becomes unstable, paving the way for chaos to emerge. In contrast to random matrices with light-tailed elements, random matrices with $\alpha$-stable entries feature an unbounded limiting density of eigenvalues (Bordenave et al., 2011). In infinite networks, this can lead to the lack of transition between quiescent and chaotic states, with any perturbation ultimately expanding in a chaotic manner (Kuśmierz et al., 2020).
@@
Understanding the computational implications of heavy-tailed recurrent connectivity is especially timely as RNNs have become central to neuroscience modeling. They are used to reproduce latent trajectories from neural recordings (Sussillo and Abbott, 2009; Rajan et al., 2016; Pandarinath et al., 2018; Keshtkaran et al., 2022), to simulate circuit mechanisms of cognitive tasks (Yang et al., 2019; Driscoll et al., 2024), and to complement experiments through hypothesis generation (Pinto et al., 2019; Pagan et al., 2025). In NeuroAI, RNNs have been embedded into deep RL agents to recapitulate biological navigation codes of grid cells (Banino et al., 2018). Yet, despite their prevalence, theoretical understanding of RNNs with biologically realistic, heavy-tailed weights remains limited.
@@
  • We derive theoretical predictions for the critical gain at which this transition occurs and validate them through simulations.
  • We show numerically that heavier-tailed RNNs exhibit a slower transition to chaos, sustaining edge-of-chaos dynamics over a broader gain regime and offering greater robustness to gain variation.
@@ 2 Related Works
Finally, a mean-field theory of Cauchy RNNs (i.e., with weights following a Lévy $\alpha$-stable distribution where $\alpha=1$), is presented in Kuśmierz et al. (2020). Specifically, they show that Cauchy RNNs with a binary activation function exhibit transition to chaos and generate scale-free avalanches, similarly observed in biological neural recordings (Beggs and Plenz, 2003) and often presented as evidence supporting the critical brain hypothesis (Muñoz, 2018). Kuśmierz et al. (2020) also note that Cauchy networks with a wide class of activation functions, including $\mathrm{tanh}$ studied in our work, are always chaotic in the infinite-size limit, and, as such, do not exhibit transition to chaos. In contrast, our results reveal that this observation is no longer true in the finite networks, highlighting the importance of finite-size effects.
@@ 3.4 Participation ratio and Lyapunov dimension
Near the edge of chaos, all positive Lyapunov exponents are close to 00 and perturbations along the corresponding directions expand with slow timescales. As a result, in this regime, a higher $D_{\mathrm{KY}}$ indicates that the system evolves on a higher-dimensional slow manifold (Krishnamurthy et al., 2022), with more modes contributing to long-term variability and slow divergence–implying a greater capacity to support rich, temporally extended computations. We track all orthogonal directions and update them with QR decomposition at each step after a fixed warmup. We examine how $D_{\mathrm{KY}}$ evolves with gain $g$ across all dynamical regimes.
@@
Let $x^{(t)}\!\in\!\mathbb{R}^{N}$ be the hidden state of the RNN at time $t$, recorded over the final $K$ steps of a length-$T$ trajectory at fixed gain $g$ with $K>N$, after discarding the initial $T-K$ warmup steps. We compute the empirical covariance matrix $S=\frac{1}{T-1}\sum_{t=1}^{T}\big{(}x^{(t)}-\bar{x}\big{)}\big{(}x^{(t)}-\bar{x}\big{)}^{\top},$ where $\bar{x}=\frac{1}{T}\sum_{t=1}^{T}x^{(t)}.$ Let $\tilde{\lambda}$ denote the eigenvalues of $S$. The participation ratio is defined as (Kramer and MacKinnon, 1993; Gao et al., 2017; Recanatesi et al., 2022):
@@ 4.1.1 Information propagation in feedforward networks
We also confirm our theoretical predictions in simulations, by passing a random initial vector through $T=100$ steps (layers) of a linearized network with weights redrawn at each step from a fixed distribution. We fix $\alpha=1$ and vary $g$. Below (above) the transition, we expect the components of the final state to be close to (far from) zero with high probability. Thus, we construct a simple order parameter $f_{<\epsilon}$ defined as the number of components of the final state $\varepsilon(T)$ that are within $\epsilon$ from 00. As shown in Fig. 1B, the network goes through a sharp transition between $f_{<\epsilon}=1$ and $f_{<\epsilon}=0$, and the location of the transition is consistent with our theoretical prediction. The transition is rather sharp even for small networks (Fig. 1C, $N=100$), and is expected to become even sharper with increasing number of steps $T$. Our theoretical result becomes exact in the limit of $T\to\infty$.
@@
Figure 1: (A): Transition point $g^{*}$ predicted by our theory as a function of network size for various $\alpha$. The transition point of Gaussian networks rapidly converges to the mean-field limit (dashed line). In contrast, the transition point of heavy-tailed networks decays slowly towards zero. (B): The fraction of small ($\epsilon=0.1$) final state components in linear networks with $\alpha=1$ and $N=3000$ evolved for $T=100$ steps from random initial conditions as a function of $g$. In the annealed case, we observe a sharp transition at the location predicted by the theory. In the quenched case, each individual realization exhibits a sharp transition (thin blue lines), but its location varies between different realizations of the weight matrix. Thus, when averaged over the realizations (thick blue line and dots; shaded region shows the ±$\pm 3$ standard error), the transition looks smoother than in the annealed case. Nonetheless, its location is approximately predicted by the theory. (C): Same as B but with $N=100$. As predicted by the theory, the transition point shifts to the right with decreasing $N$. Moreover, the location of the transition in the quenched case varies more in smaller networks.
@@ 4.1.2 Quenched disorder and recurrent neural networks
To test this hypothesis, we simulate a quenched version of (6) in which $\forall_{t}W^{(t)}=W$. In order to observe the transition, we first fix the random seed (i.e., draw random components of the weight matrix from $L_{\alpha}(1/N^{1/\alpha})$) and then rescale its components by various values of $g$. We focus on $\alpha=1$ as a representative example. As shown in Fig. 1B, evolution of each realization of the weight matrix goes through a very sharp transition, but the location of this transition varies significantly between the realizations. Nonetheless, they are concentrated around the point predicted by the annealed theory. Moreover, the location of the transition shifts to the right and fluctuations increase with decreasing $N$ and (Fig. 1C). These results suggest that the location of the transition in the quenched case approaches the annealed prediction with increasing $N$.
@@ 4.2 Heavier-tailed RNNs exhibit a slower, more robust transition to chaos
In networks with Gaussian connectivity ($\alpha=2.0$), the maximal Lyapunov exponent (MLE) increases steeply with gain $g$, indicating a rapid onset of chaos. In contrast, RNNs with heavier-tailed weights (lower $\alpha$) exhibit a slower rise in the MLE as $g$ increases near the transition (when the MLE is near zero). This gradual transition implies that these networks remain closer to the edge of chaos over a wider range of gain values, consistent with previous observations of an extended, critical-like region (Wardak and Gong, 2022; Qu et al., 2022). Such extended critical-like behavior can offer a form of robustness with respect to changes in network parameters, which can be an important property that benefits biological networks in non-stationary environments, allowing the network to maintain sensitive high-capacity dynamics (Bertschinger et al., 2004; Legenstein and Maass, 2007; Toyoizumi and Abbott, 2011) without the requirement of precise parameter adjustment.
@@
Figure 2: Maximum Lyapunov exponent ($\lambda_{\text{max}}$) as a function of gain $g$ for autonomous recurrent networks with different tail indices $\alpha$, shown for: (A) $N=1000$, (B) $N=3000$, and (C) $N=10000$. Curves show mean across 10 trials; shaded regions denote $\pm$1 SD. We let the networks evolve for $T=3000$ steps, among which the Lyapunov exponents are accumulated over the last $K=100$ steps. See results under noisy stimulus and ablation studies in Appendices E, G. Heavier-tailed networks (lower $\alpha$) exhibit a slower, more gradual increase in $\lambda_{\text{max}}$ near the transition (where $\lambda_{\text{max}}=0$), resulting in a broader edge-of-chaos regime with respect to $g$. Dashed lines and legend mark the average critical gain $g^{*}$ at which $\lambda_{\text{max}}$ first crosses zero. As $N$ increases, this transition shifts leftward, especially for lower $\alpha$, in line with our theoretical predictions on finite-size effects.

### v2

@@ Abstract
Growing evidence suggests that synaptic weights in the brain follow heavy-tailed distributions, yet most theoretical analyses of recurrent neural networks (RNNs) assume Gaussian connectivity. We systematically study the activity of RNNs with random weights drawn from biologically plausible Lévy alpha-stable distributions. While mean-field theory for the infinite system predicts that the quiescent state is always unstable—implying ubiquitous chaos—our finite-size analysis reveals a sharp transition between quiescent and chaotic dynamics. We theoretically predict the gain at which the finite system transitions from quiescent to chaotic dynamics, and validate it through simulations. Compared to Gaussian networks, finite heavy-tailed RNNs exhibit a broader gain regime near the edge of chaos, namely, a slow transition to chaos. However, this robustness comes with a tradeoff: heavier tails reduce the Lyapunov dimension of the attractor, indicating lower effective dimensionality. Our results reveal a biologically aligned tradeoff between the robustness of dynamics near the edge of chaos and the richness of high-dimensional neural activity. By analytically characterizing the transition point in finite-size networks—where mean-field theory breaks down—we provide a tractable framework for understanding dynamics in realistically sized, heavy-tailed neural circuits.***The codebase is publicly available at https://github.com/AllenInstitute/HeavyRNN_public.
@@ 1 Introduction
Advances in connectomics yield increasingly detailed wiring diagrams of neural circuits across species and brain regions (Dorkenwald et al., 2024; The MICrONS Consortium, 2025). This progress raises fundamental questions: what structural principles govern neural circuits, and how do they support the brain’s remarkable computational power? One such prominent structural feature is the presence of heavy-tailed (Foss et al., 2011) synaptic weight distributions, consistently observed across the mammalian cortex (Song et al., 2005; Lefort et al., 2009; Dorkenwald et al., 2022), mammalian hippocampus (Ikegaya et al., 2013), and even in the Drosophila central brain (Scheffer et al., 2020). Notably, this feature stands in sharp contrast to the Gaussian weight assumptions that dominate theoretical neuroscience studies and light-tailed distributions utilized in the standard initialization schemes in modern artificial neural networks (LeCun et al., 2002; Glorot and Bengio, 2010; He et al., 2015). One way to formally model heavy tails is with the family of Lévy $\alpha$-stable distributions (Feller, 1971; Borak et al., 2005), which emerges as a natural generalization of the familiar Gaussian distribution ($\alpha=2$) via the generalized central limit theorem. The family is parameterized by a stability index $\alpha$, where smaller values of $\alpha$ correspond to heavier tails. For $\alpha<2$, these distributions feature heavy, power-law tails. Similarly to experimentally measured synaptic weights, samples generated from such distributions consistently contain large outliers that dominate many sample statistics. As we show in this paper, this can strongly affect neural dynamics.
One key phenomenon studied in theoretical neuroscience and machine learning is the transition to chaos, which has long been hypothesized to support optimal information flow and computational capacity at the so-called edge of chaos in a wide variety of randomly initialized neural networks. This encompasses RNNs (Bertschinger et al., 2004; Legenstein and Maass, 2007; Toyoizumi and Abbott, 2011; Schuecker et al., 2018) and feedforward neural networks (Schoenholz et al., 2016; Poole et al., 2016). In particular, Schoenholz et al. (2016) shows that the trainability of deep networks depends on initializing near the edge of chaos: the farther from this critical regime, the shallower a network must be to remain trainable. Similarly, Bertschinger et al. (2004) shows that only near the edge of chaos can RNNs perform complex computations on time series.
@@
In the context of feedforward networks, this effect can be understood in terms of the neural network Gaussian process kernel (Neal, 1996): outside of the critical point, analogous to the edge of chaos in RNNs, a trivial fixed point with kernel constant almost everywhere is approached exponentially fast, limiting the effective depth of information propagation and network trainability (Schoenholz et al., 2016; Lee et al., 2017). The transition between chaotic and non-chaotic dynamics in RNNs is often discussed in terms of eigenvalues of the weight matrix (Rajan and Abbott, 2006; Aljadeff et al., 2015). According to the circular law (Girko, 1985; Tao et al., 2010), its eigenvalues are bounded in a circle of radius proportional to the standard deviation of the distribution of weight entries. If the standard deviation is large enough, some eigenvalues fall outside of the unit circle and the quiescent state becomes unstable, paving the way for chaos to emerge. In contrast to random matrices with light-tailed elements, random matrices with $\alpha$-stable entries feature an unbounded limiting density of eigenvalues (Bordenave et al., 2011). In infinite networks, this can lead to the lack of transition between quiescent and chaotic states, with any perturbation ultimately expanding in a chaotic manner (Kuśmierz et al., 2020).
@@
Understanding the computational implications of heavy-tailed recurrent connectivity is especially timely as RNNs have become central to neuroscience modeling. They are used to reproduce latent trajectories from neural recordings (Sussillo and Abbott, 2009; Rajan et al., 2016; Pandarinath et al., 2018; Keshtkaran et al., 2022), to simulate circuit mechanisms of cognitive tasks (Yang et al., 2019; Driscoll et al., 2024), and to complement experiments through hypothesis generation (Pinto et al., 2019; Pagan et al., 2025). In NeuroAI, RNNs have been embedded into deep reinforcement learning agents to recapitulate biological navigation codes of grid cells (Banino et al., 2018). Yet, despite their prevalence, theoretical understanding of RNNs with biologically realistic, heavy-tailed weights remains limited.
@@
  • We derive theoretical predictions for the critical gain at which this transition occurs as a function of network size, and validate them through simulations.
  • We show numerically that heavier-tailed RNNs exhibit a slower transition to chaos, sustaining edge-of-chaos dynamics over a broader gain regime and offering greater robustness to gain variation; we show this can translate to improved information processing, as evidenced by the superior performance of heavy-tailed RNNs on a simple reservoir-computing task.
@@ 2 Related Works
Additionally, a mean-field theory of Cauchy RNNs (i.e., with weights following a Lévy $\alpha$-stable distribution where $\alpha=1$), is presented in Kuśmierz et al. (2020). Specifically, they show that Cauchy RNNs with a binary activation function exhibit transition to chaos and generate scale-free avalanches, similarly observed in biological neural recordings (Beggs and Plenz, 2003) and often presented as evidence supporting the critical brain hypothesis (Muñoz, 2018). Notably, Kuśmierz et al. (2020) note that Cauchy networks with a wide class of activation functions, including $\mathrm{tanh}$ studied in our work, are always chaotic in the infinite-size limit, and, as such, do not exhibit a transition to chaos. In contrast, our results reveal that this observation is no longer true in the finite networks, highlighting the importance of finite-size effects.
Finally, our work complements recent studies on brain-like learning with exponentiated gradients (Cornford et al., 2024), which showed that such updates naturally give rise to log-normal connectivity distributions. Within this broader context, our results offer a theoretical perspective that elucidates the dynamical consequences of these heavy-tailed structures.
@@ 3.4 Participation ratio and Lyapunov dimension
Near the edge of chaos, all positive Lyapunov exponents are close to 0 and perturbations along the corresponding directions expand with slow timescales. As a result, in this regime, a higher $D_{\mathrm{KY}}$ indicates that the system evolves on a higher-dimensional slow manifold (Krishnamurthy et al., 2022), with more modes contributing to long-term variability and slow divergence–implying a greater capacity to support rich, temporally extended computations. We track all orthogonal directions and update them with QR decomposition at each step after a fixed warmup. We examine how $D_{\mathrm{KY}}$ evolves with gain $g$ across all dynamical regimes.
@@
Let $x^{(t)}\!\in\!\mathbb{R}^{N}$ be the hidden state of the RNN at time $t$, recorded over the final $K$ steps of a length-$T$ trajectory at fixed gain $g$ with $K>N$, after discarding the initial TKT-K warmup steps. We compute the empirical covariance matrix $S=\frac{1}{T-1}\sum_{t=1}^{T}\big(x^{(t)}-\bar{x}\big)\big(x^{(t)}-\bar{x}\big)^{\top},$ where $\bar{x}=\frac{1}{T}\sum_{t=1}^{T}x^{(t)}.$ Let $\tilde{\lambda}$ denote the eigenvalues of $S$. The participation ratio is defined as (Kramer and MacKinnon, 1993; Gao et al., 2017; Recanatesi et al., 2022):
@@ 4.1.1 Information propagation in feedforward networks
We also confirm our theoretical predictions in simulations, by passing a random initial vector through $T=100$ steps (layers) of a linearized network with weights redrawn at each step from a fixed distribution. We fix $\alpha=1$ and vary $g$. Below (above) the transition, we expect the components of the final state to be close to (far from) zero with high probability. Thus, we construct a simple order parameter $f_{<\epsilon}$ defined as the number of components of the final state $\varepsilon(T)$ that are within $\epsilon$ from 0. As shown in Fig. 1B, the network goes through a sharp transition between $f_{<\epsilon}=1$ and $f_{<\epsilon}=0$, and the location of the transition is consistent with our theoretical prediction. The transition is rather sharp even for small networks (Fig. 1C, $N=100$), and is expected to become even sharper with increasing number of steps $T$. Our theoretical result becomes exact in the limit of $T\to\infty$.
@@
Figure 1: (A): Transition point $g^{*}$ predicted by our theory as a function of network size for various $\alpha$. The transition point of Gaussian networks rapidly converges to the mean-field limit (dashed line). In contrast, the transition point of heavy-tailed networks decays slowly towards zero. (B): The fraction of small ($\epsilon=0.1$) final state components in linear networks with $\alpha=1$ and $N=3000$ evolved for $T=100$ steps from random initial conditions as a function of $g$. In the annealed case, we observe a sharp transition at the location predicted by the theory. In the quenched case, each individual realization exhibits a sharp transition (thin blue lines), but its location varies between different realizations of the weight matrix. Thus, when averaged over the realizations (thick blue line and dots; shaded region shows the $\pm 3$ standard error), the transition looks smoother than in the annealed case. Nonetheless, its location is approximately predicted by the theory. (C): Same as B but with $N=100$. As predicted by the theory, the transition point shifts to the right with decreasing NN. Moreover, the location of the transition in the quenched case varies more in smaller networks.
@@ 4.1.2 Quenched disorder and recurrent neural networks
To test this hypothesis, we simulate a quenched version of (6) in which $\forall_{t}W^{(t)}=W$. In order to observe the transition, we first fix the random seed (i.e., draw random components of the weight matrix from $L_{\alpha}(1/N^{1/\alpha})$) and then rescale its components by various values of $g$. We focus on $\alpha=1$ as a representative example. As shown in Fig. 1B, evolution of each realization of the weight matrix goes through a very sharp transition, but the location of this transition varies significantly between the realizations. Nonetheless, they are concentrated around the point predicted by the annealed theory. Moreover, the location of the transition shifts to the right and fluctuations increase with decreasing $N$ and (Fig. 1C). These results suggest that the location of the transition in the quenched case approaches the annealed prediction with increasing $N$. We provide further analysis on the behavior of the quanched transition point as a function of network size $N$ in Appendix J.
@@
While our analysis focuses on the tanh activation, it generalizes to any function satisfying $\phi(0)=0$ and admitting a local expansion $\phi(x)=ax+o(x)$. In this regime, the existence of the transition follows directly from the linear stability of the quiescent fixed point. For unbounded activations such as ReLU, however, bounded dynamics are no longer guaranteed, and divergence may occur at large $g$. Although the transition itself persists for a broad class of activations, the qualitative behavior above it can differ substantially. Beyond the transition, the ensuing dynamics depend sensitively on the nonlinearity: linear or ReLU activations typically diverge for large $g$, whereas sublinear, saturating nonlinearities constrain activity and preserve stability. We therefore expect our results to hold for any smooth, saturating activation function, while unbounded ones likely produce more complex, divergent dynamics. The framework can also be extended to cases with $\phi(0)\neq 0$ by expanding around the corresponding non-quiescent fixed point. Here, the fixed point’s location may vary with the order parameter $g$, but the overall nature of the transition should remain unchanged.
@@ 4.2 Heavier-tailed RNNs exhibit a slower, more robust transition to chaos
In networks with Gaussian connectivity ($\alpha=2.0$), the maximal Lyapunov exponent (MLE) increases steeply with gain $g$, indicating a rapid onset of chaos. In contrast, RNNs with heavier-tailed weights (lower $\alpha$) exhibit a slower rise in the MLE as $g$ increases near the transition (when the MLE is near zero). This gradual transition implies that these networks remain closer to the edge of chaos over a wider range of gain values, consistent with previous observations of an extended, critical-like region (Wardak and Gong, 2022; Qu et al., 2022). Such extended critical-like behavior can offer a form of robustness with respect to changes in network parameters, which can be an important property that benefits biological networks in non-stationary environments, allowing the network to maintain sensitive high-capacity dynamics (Bertschinger et al., 2004; Legenstein and Maass, 2007; Toyoizumi and Abbott, 2011) without the requirement of precise parameter adjustment. In our analysis of reservoir-computing networks on a delayed XOR task (Appendix L), heavy-tailed networks maintained strong task performance across a broader gain regime than Gaussian networks. This provides a concrete proof-of-concept that the extended critical regime enhances robustness and performance without fine-tuning, which may benefit both machine learning applications and neural computation. Future studies could extend this analysis to trained recurrent networks and more complex temporal tasks to further elucidate how heavy-tailed connectivity shapes information processing and learning dynamics.
@@
Figure 2: Maximum Lyapunov exponent ($\lambda_{\text{max}}$) as a function of gain $g$ for autonomous recurrent networks with different tail indices $\alpha$, shown for: (A) $N=1000$, (B) $N=3000$, and (C) $N=10000$. Curves show mean across 10 trials; shaded regions denote $\pm$1 SD. We let the networks evolve for $T=3000$ steps, among which the Lyapunov exponents are accumulated over the last $K=100$ steps. See results under noisy stimulus and ablation studies in Appendices E, G. Heavier-tailed networks (lower $\alpha$) exhibit a slower, more gradual increase in $\lambda_{\text{max}}$ near the transition (where $\lambda_{\text{max}}=0$), resulting in a broader edge-of-chaos regime with respect to $g$. Dashed lines and legend mark the average critical gain $g^{*}$ at which $\lambda_{\text{max}}$ first crosses zero. As NN increases, this transition shifts leftward, especially for lower $\alpha$, in line with our theoretical predictions on finite-size effects.